Dry steam at 100 0c is bubbled into 500g of water originally at 200c. What will be the temperature of the water after 30g of steam has condensed?
1
Expert's answer
2018-02-14T08:26:07-0500
mL+mc(T_s-T)=Mc(T-T_w )
30(2265)+30(4.18)(100-T)=500(4.18)(T-20)
T=55℃.